OpenStudy (anonymous):

A force F=(3,5,12) in newtons is applied to lift a box with displacement S=(0,0,8). Calculate the work against gravity and the work in the direction of travel.

OpenStudy (anonymous):

(x,y,z) I'm completely sure on this But it appears the box was only displaced in the z direction. So the only work done is 12*8=96. work done against gravity..0?

OpenStudy (anonymous):

what if it were displaced (2,0,10)?

OpenStudy (anonymous):

@cherio12

OpenStudy (anonymous):

okay so now you've moved the object in both the x and z direction. Work in the x direction = 2*3=6 Work in the z direction = 10*12=120

OpenStudy (anonymous):

assuming i'm doing this correctly..

OpenStudy (anonymous):

that's right, now what about the work against gravity?

OpenStudy (anonymous):

Well the net work in the y direction is 0 since the object doesn't displace this way. I know it says that a force of 5 is applied in the y direction..but i don't quite know where that goes. perhaps the work against gravity is 5...but that shouldnt be..

OpenStudy (anonymous):

would i only take the z component into consideration?

OpenStudy (anonymous):

for the first one?...sorry ive never really seen force slash displacement in this form before